# Deep Unfolding with Approximated Computations for Rapid Optimization

### v1

@@ IV Case Study 1: Hybrid Beamforming
This section presents an application of our methodology within an unsupervised learning setup, as introduced in Subsection III-B3, focusing on the task of hybrid beamforming [34], which is considered to be key in enabling high-frequency large-scale multiple-input multiple-output (MIMO) systems [35]. Our motivation for using hybrid beamforming as a case study stems not only from its technological importance, but also from its need for rapid optimization solvers, since: $(i)$ The beamforming task is typically represented as an optimization problem that is tackled with iterative solvers [36, 37, 38, 39]; $(ii)$ Hybrid beamformers are tuned for a given channel state information (CSI), which can change rapidly (over $10^{3}$ times a second [3]); $(iii)$ While unfolded optimizers have been shown to yield suitable hybrid beamformers within a few iterations [21, 22, 23, 40, 41], each iteration is often computationally intense, especially in wideband regimes, such that the optimizer cannot be applied within a coherence duration. Accordingly, this case study represents an optimization setup where there is a concrete need to simultaneously reduce the number of iterations and alleviate the burden within the iterations.
@@ IV-A1 Channel Model
We consider a single-cell downlink MIMO system comprising $N$ users and a base station (BS) with $M$ antennas. The BS utilizes $B$ frequency bands shared among all users in a non-orthogonal fashion. We let ${\bm{c}}_{b}\in~\mathbb{C}^{N}$ denote symbols transmitted at the $b$th frequency bin, $b\in\{1,2,\ldots,B\}\triangleq\mathcal{B}$, The symbols are i.i.d and have equal power such that $\mathbb{E}[{\bm{c}}_{b}{\bm{c}}_{b}^{H}]=\frac{1}{N}\textbf{I}_{N}$ for each $b\in\mathcal{B}$.
The outgoing ${\bm{c}}_{b}$ is precoded into the channel input ${\bm{p}}_{b}\in\mathbb{C}^{M}$, and corresponding channel outputs can be written as
@@
Here, ${\bm{H}}_{b}\in\mathbb{C}^{N\times M}$ is the channel at the $b$th band, and ${\bm{n}}_{b}\in\mathbb{C}^{N}$ is Gaussian noise with i.i.d. entries of variance $\sigma^{2}$.
@@ IV-A2 Hybrid Beamforming
The BS has $L<M$ RF chains. The symbols $\{{\bm{c}}_{b}\}_{b\in\mathcal{B}}$ are thus precoded in two stages. First, a digital precoder ${\bm{W}}_{d,b}\in\mathbb{C}^{L\times{N}}$ is applied to ${\bm{c}}_{b}$ in each frequency bin $b\in\mathcal{B}$. Next, the digital symbols are combined into the channel input ${\bm{p}}_{b}$ using an analog precoder. Analog precoding is carried out using phase shifters which are static in frequency [34], and is thus modeled via the matrix ${\bm{W}}_{a}\in\mathcal{A}\subseteq\mathbb{C}^{M\times{L}}$, where $\mathcal{A}$ represents the set of matrices where the entries have a unit magnitude, namely
@@
To conclude, the channel input after precoding is ${\bm{p}}_{b}={\bm{W}}_{a}{\bm{W}}_{d,b}\cdot{\bm{c}}_{b}$ for bin $b$. Thus, according to (12), the overall channel input-output relationship is
${\bm{y}}_{b}={\bm{H}}_{b}{\bm{W}}_{a}{\bm{W}}_{d,b}\cdot{\bm{c}}_{b}+{\bm{n}}_{b},\qquad\forall b\in\mathcal{B}.$ (14)
The precoder satisfies the following power constraint
$\frac{1}{B}\sum\nolimits_{j=1}^{B}\|{\bm{W}}_{a}{\bm{W}}_{d,j}\|^{2}_{F}\leq\emph{N}.$ (15)
Figure 1: Hybrid beamforming case study illustration
@@ IV-A3 Problem Formulation
Hybrid beamforming design refers to setting the precoders ${\bm{W}}_{a},\{{\bm{W}}_{d,b}\}$ for a given CSI $\{{\bm{H}}_{b}\}$. Namely, the decision here is ${\bm{s}}=[{\bm{W}}_{a},\{{\bm{W}}_{d,b}\}]$, and the context is ${\bm{x}}=\{{\bm{H}}_{b}\}$, as illustrated in Fig. 1. We focus on sum-rate maximization, being a common measure for the quality of a resulting channel that is invariant of receiver-side processing [42, 43, 44]. The BS has accurate CSI, i.e., knowledge of $\{{\bm{H}}_{b}\}$, and thus the following sum-rate is achievable [45]
@@
Accordingly, the hybrid beamforming task can be written as tackling the following optimization problem
@@ IV-B Iterative Optimizer
Problem (17) is a constrained maximization, which can be tackled using PGA with alternating optimization [21]. In this method, each iteration first optimizes ${\bm{W}}_{a}$ while keeping $\{{\bm{W}}_{d,b}\}$ fixed, then repeats this process for every ${\bm{W}}_{d,b}$. Each matrix update is followed by a projection to ensure that ${\bm{W}}_{a}\in\mathcal{A}$ and the power constraint in (15) is satisfied.
Specifically, the $k$th iteration updates ${\bm{W}}_{a}$ via
@@ IV-C Learned Approximated Optimizer
While PGA can yield suitable hybrid precoders, it is often unsuitable for real-time operation (i.e. within a coherence duration) since $(i)$ it requires multiple iterations to converge; and $(ii)$ each iteration involves heavy computations. Specifically, obtaining the gradients in (19), carried out in each iteration, has complexity of [21]
@@
Applying our methodology outlined in Subsection III-B for tackling (17) with PGA leads to Learned Approximated (LAPGA), which occasionally replaces gradient computations with low-complexity approximations, leveraging data and increased parameterization to maintain suitable precoders and preserve the operation of PGA. Accordingly, we first fix the number of iterations to a small integer $K$. To also reduce the per-iteration complexity, we introduce the following approximations:
@@
    $\tilde{{R}}_{b}^{(k)}=\begin{cases}{R}({\bm{W}}_{a}^{(k+1)},\{{\bm{W}}_{d,\tilde{b}}^{(k)}\},\{{\bm{H}}_{\tilde{b}}\})&k\in\mathcal{K}_{d,b}^{\rm approx},\\ {R}({\bm{W}}_{a}^{(k)},\{{\bm{W}}_{d,\tilde{b}}^{(k-1)}\},\{{\bm{H}}_{\tilde{b}}\})&k\notin\mathcal{K}_{d,b}^{\rm approx}.\end{cases}$
@@
Combining I1-I3, the resulting LAPGA (using the initiation proposed in [21]) is summarized as Algorithm 2. LAPGA is trained using data $\mathcal{D}$ comprised of past channel realizations, written as $\mathcal{D}=\{\{\tilde{{\bm{H}}}_{b}^{r}\}_{b\in\mathcal{B}}\}_{r=1}^{|\mathcal{D}|}$, i.e., there is no "ground-truth" precoders. Accordingly, training is carried out based on the unsupervised empirical risk (10), using, e.g., Algorithm 1.
Init : Step sizes $\{\bm{\mu}_{d,b}^{(k)}\}_{b\in\mathcal{B}},\bm{\mu}_{a}^{(k)}$; $\{{\bm{W}}_{d,b}^{(0)}\}_{b\in\mathcal{B}}$;
Approximation indices $\mathcal{K}_{a}^{\rm approx}$, $\{\mathcal{K}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}d,}b}\}_{b\in\mathcal{B}}$;
@@
6    ${\bm{W}}_{d,b}^{(k\!+\!1)}\!\leftarrow\Pi_{\mathcal{P}}\left\{{\bm{W}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a}}^{(k)}\!+\!\bm{\mu}_{d,b}^{(k)}\odot\frac{\partial{\tilde{R}_{b}^{(k)}}}{\partial{\bm{W}}_{d,b}}\right\}$;
@@
LAPGA realizes an iterative optimizer with few iterations and approximated low-complexity operations, while leveraging data to mitigate performance reduction. In particular, the average complexity of each iteration is given by
@@
which follows since the complexity in (20) is obtained by equal contributions of the digital and analog settings [21]. The reduction in (22) is shown in Section IV-D to have a significant impact in latency, with minimal impact in performance. The resulting insight, i.e., that deep unfolding can reduce both iterations and the complexity of each iteration, is expected to be relevant beyond our considered domain of PGA-based hybrid precoding.
IV-D Experimental Study
In this section, we numerically evaluate LAPGA. Our primary objective is to demonstrate that our proposed method achieves comparable results to non-approximated solvers, while significantly reducing online computations. We thus use the following benchmarks, selected due to their relationship with LAPGA, while being suitable algorithms for hybrid beamforming compared to alternative methods [3].
@@ IV-D Experimental Study
    Unfolded full PGA with scalar step-sizes of [21].
@@
    Classic PGA with fixed step-sizes optimized manually.
We first consider a small-scale MIMO channel, with $B=8$ , $N=6$, $L=10$ and $M=12$. The channels are obtained from the QuadRiGa model [46], using 1000 channels for training and 100 channels for test. We implement LAPGA employing I1-I3, while setting $\mathcal{K}_{a}^{\rm approx}=\{1,\ldots,K\}$ and $\mathcal{K}_{d,b}^{\rm approx}$ being even indices. We use this setup to also assess the effect of its approximations. We compare LAPGA to B1-B2 as well as its following variations:
@@
Fig. 2 reports the average per-user sum-rate after $K$ iterations (set to $K=5$ except for B2, that uses $K=50$) versus signal-to-noise ratio (SNR), defined as $1/N\sigma^{2}$. We observe in Fig. 2 that unfolding can lead to improved performance. Specifically, B1 achieves better results than B2 while requiring $66\%-90\%$ less iterations, while LAPGA surpasses the achievable rate obtained by B1 while requiring $70\%$ less calculations. Overall, our proposed algorithm outperforms B2 while requiring between $89\%-97.3\%$ less calculations, depending on the SNR. To see this, we report in Table I the average and standard deviation of the rate achieved at SNR of 0 dB, along with the overall number of products used to map the CSI to the precoders. Table I reveals the dramatic complexity reduction of LAPGA. It also shows that employing approximations I1-I2 allows achieving different balances between performance and complexity, and that the increased parameterization in I3 greatly facilitates learning to cope with their induced errors.
@@
We proceed to a larger MIMO setting, with $B=64$ , $N=12$, $L=12$ and $M=32$. Having assessed the individual contributions of I1-I3 in the previous study, here we consider LAPGA employing all I1-I3. We again set $K=5$ for the unfolded optimizers, while B2 uses $K=100$ iterations.
@@
The resulting rate vs. SNR is reported in Fig. 3, while Table II also includes complexity at 0 dB SNR. As can be seen, LAPGA achieves better results in both high and low SNR scenarios compared to B1 and B2. In the mid-range SNR, B1 performs slightly better, but requires 70% more computations to do so, with B2 being three orders of magnitude more complex. These results indicate the potential of our unfolded approximated framework to greatly facilitate low-complexity rapid optimization.
@@ V Case Study 2: Robust PCA
The previous case study demonstrated the utility of our proposed methodology in designing rapid optimization solvers in a domain where the optimization objective is both practically significant and computationally intensive. Through the hybrid beamforming study, we highlighted how unfolding combined with approximated computations can enable efficient inference with drastically reduced latency. In this section, we explore the application of RPCA, which is a widely-used task in various fields ranging from background subtraction and anomaly detection in image and video processing [47]. RPCA serves as a representative setup for optimization problems that are $(i)$ computationally demanding and have motivated several unfolded, yet non-approximated, optimization strategies in prior work [48, 26, 27]; $(ii)$ formulated using surrogate mathematical objectives rather than the true task loss. In particular, conventional RPCA formulations rely on convex relaxations of the underlying objective [29], which may be misaligned with the ultimate performance metric of interest. This makes RPCA an especially suitable domain for examining the added value of our approach in scenarios where the learning objective deviates from the optimization one.
To describe this case study, we first formulate RPCA and the considered first-order optimizer in Subsections V-A-V-B, respectively. Then we detail our learned unfolded methodology in Subsection V-C, and present the associated experimental evaluation in Subsection V-D.

### v2

@@ IV Case Study 1: Hybrid Beamforming
This section presents an application of our methodology within an unsupervised learning setup, as introduced in Subsection III-B3, focusing on the task of hybrid beamforming [3], which is considered key in enabling high-frequency large-scale multiple-input multiple-output (MIMO) systems [35]. Our motivation for using hybrid beamforming as a case study stems not only from its technological importance, but also from its need for rapid optimization solvers, since: $(i)$ The beamforming task is typically represented as an optimization problem that is tackled with iterative solvers [36, 37, 38, 39]; $(ii)$ Hybrid beamformers are tuned for a given channel state information (CSI), which can change rapidly (over $10^{3}$ times a second [3]); $(iii)$ While unfolded optimizers have been shown to yield suitable hybrid beamformers within a few iterations [22, 23, 24, 40, 41], each iteration is often computationally intense, especially in wideband regimes, such that the optimizer cannot be applied within a coherence duration. Accordingly, this case study represents an optimization setup where there is a concrete need to simultaneously reduce the number of iterations and alleviate the burden within the iterations. While this case study focuses on alternating optimization based on PGA, selected as a representative baseline due to its widespread use, its established empirical performance, and its suitability for deep unfolding [22], the proposed methodology is generic and can be applied to alternative hybrid beamforming optimizers as well. Our emphasis is not on the specific choice of solver, but on demonstrating how iterative algorithms can be transformed into reduced-complexity, trainable models via approximated deep unfolding.
@@ IV-A1 Channel Model
Consider a single-cell downlink MIMO system featuring an $M$-antenna base station (BS) and $N$ users. The BS utilizes $B$ frequency bands, shared among all users in a non-orthogonal fashion. Let $\mathbf{s}_{b}\in\mathbb{C}^{N}$ denote the symbol vector transmitted at the $b$-th frequency bin, where $b\in\mathcal{B}\triangleq\{1,2,\dots,B\}$. The symbols are assumed to be i.i.d. with equal power such that $\mathbb{E}[\mathbf{s}_{b}\mathbf{s}_{b}^{H}]=\frac{1}{N}\mathbf{I}_{N}$ for each $b\in\mathcal{B}$.
We precode the symbols $\mathbf{s}_{b}$ into the channel input ${\bm{p}}_{b}\in\mathbb{C}^{M}$, yielding the channel output:
@@
Where ${\bm{H}}_{b}\in\mathbb{C}^{N\times M}$ is the channel at the $b$th band, and ${\bm{n}}_{b}\in\mathbb{C}^{N}$ is Gaussian noise with i.i.d. entries of variance $\sigma^{2}$.
@@ IV-A2 Hybrid Beamforming
With $L<M$ Radio Frequency (RF) chains at the BS, a two-stage hybrid precoding scheme is employed for the symbols $\{\mathbf{s}_{b}\}_{b\in\mathcal{B}}$. In the first stage, a frequency-dependent digital precoder $\mathbf{W}_{d,b}\in\mathbb{C}^{L\times N}$ is applied. In the second stage, an analog precoder $\mathbf{W}_{a}\in\mathcal{A}$ combines these signals into the channel input ${\bm{p}}_{b}$. Following [3], $\mathbf{W}_{a}$ is frequency-invariant and consists of phase shifters restricted to the unit-magnitude set $\mathcal{A}$, namely:
@@
The precoded channel input for the $b$-th frequency bin is thus
${\bm{p}}_{b}=\mathbf{W}_{a}\mathbf{W}_{d,b}\mathbf{s}_{b}.$ (14)
Based on (12), the channel input-output relationship is
$\mathbf{y}_{b}=\mathbf{H}_{b}\mathbf{W}_{a}\mathbf{W}_{d,b}\mathbf{s}_{b}+\mathbf{n}_{b},\quad\forall b\in\mathcal{B}.$ (15)
Considering the normalized symbol power, the hybrid precoders are subject to an average power constraint across all frequency bins, which is formulated as
$\frac{1}{B}\sum_{j=1}^{B}\|\mathbf{W}_{a}\mathbf{W}_{d,j}\|^{2}_{F}\leq N.$ (16)
@@ IV-A3 Problem Formulation
The hybrid beamforming design objective is to determine the precoders $\mathbf{W}_{a}$ and $\{\mathbf{W}_{d,b}\}$ for a given set of CSI $\{\mathbf{H}_{b}\}$. Namely, the decision here is ${\bm{s}}=[{\bm{W}}_{a},\{{\bm{W}}_{d,b}\}]$, and the context is ${\bm{x}}=\{{\bm{H}}_{b}\}$, as illustrated in Fig. 1. Due to its unsupervised setting, we utilize the objective function of the original optimization problem as our empirical risk as discussed in Section III-B3. Specifically, we focus on maximizing the sum-rate, a standard performance metric that is invariant of receiver-side processing. Assuming the BS possesses perfect CSI, the achievable sum-rate is given by [42]:
@@
Consequently, the task of determining the hybrid precoders can be cast as the following optimization problem:
@@ IV-B Iterative Optimizer
The constrained maximization problem (18) can be addressed, among other techniques, using PGA combined with alternating optimization [22]. In this framework, each iteration alternates between optimizing the analog precoder $\mathbf{W}_{a}$ with fixed $\{\mathbf{W}_{d,b}\}$, and subsequently updating the digital precoders while $\mathbf{W}_{a}$ remains constant. Each update is followed by a projection step to ensure that $\mathbf{W}_{a}\in\mathcal{A}$ and the power constraint in (16) are satisfied.
Specifically, the $k$-th iteration updates $\mathbf{W}_{a}$ via
@@
Figure 1: Hybrid beamforming case study illustration
@@ IV-C Learned Approximated Optimizer
While PGA can yield suitable hybrid precoders, it is often impractical for real-time operation within a channel coherence duration. Specifically, the Hybrid Beamforming task is subject to the core limitations of classical iterative methods discussed in Section IV-A3; it requires a large number of iterations to converge (C1), and suffers from high per-iteration complexity (C3) due to gradient calculations. Following [22], the computational cost per iteration is given by
@@
Following the methodology outlined in Subsection III-B, we propose Learned Approximated (LAPGA) to solve (18). LAPGA leverages data-driven parameterization to replace standard gradient computations with low-complexity approximations, thereby maintaining precoder performance while preserving the iterative structure of PGA. Specifically, we address C1 by fixing the number of iterations to a small integer $K$. Furthermore, to mitigate C3, we reduce the per-iteration complexity by employing the following approximations:
@@
    $\tilde{{R}}_{b}^{(k)}=\begin{cases}{R}({\bm{W}}_{a}^{(k+1)},\{{\bm{W}}_{d,\tilde{b}}^{(k)}\},\{{\bm{H}}_{\tilde{b}}\})&k\notin\mathcal{K}_{d,b}^{\rm approx},\\ {R}({\bm{W}}_{a}^{(k)},\{{\bm{W}}_{d,\tilde{b}}^{(k-1)}\},\{{\bm{H}}_{\tilde{b}}\})&k\in\mathcal{K}_{d,b}^{\rm approx}.\end{cases}$
@@
We note that numerous combinations of approximations, as in I1-I2, are possible. The specific configuration in LAPGA was selected as it represents a robust operating point for the complexity-performance balance. This choice leverages the unit-magnitude constraints of the analog array and the temporal stability of digital precoding iterations to maintain sum-rate while minimizing computational overhead. However, the approximation setup can be tailored to meet the specific requirements of different deployment scenarios. Combining I1-I3, the resulting LAPGA (using the initiation proposed in [22]) is summarized as Algorithm 2. LAPGA is trained using data $\mathcal{D}$ comprised of past channel realizations, written as $\mathcal{D}=\{\{\tilde{{\bm{H}}}_{b}^{r}\}_{b\in\mathcal{B}}\}_{r=1}^{|\mathcal{D}|}$, i.e., there is no "ground-truth" precoders. Accordingly, training is carried out based on the unsupervised empirical risk (10), using, e.g., Algorithm 1.
Init : Step-sizes $\{\bm{\mu}_{d,b}^{(k)}\}_{b\in\mathcal{B}},\bm{\mu}_{a}^{(k)}$; $\{{\bm{W}}_{d,b}^{(0)}\}_{b\in\mathcal{B}}$;
Approximation indices $\mathcal{K}_{a}^{\rm approx}$, $\{\mathcal{K}_{{d,}b}^{\rm approx}\}_{b\in\mathcal{B}}$;
@@
6    ${\bm{W}}_{d,b}^{(k\!+\!1)}\!\leftarrow\Pi_{\mathcal{P}}\left\{{\bm{W}}_{{a}}^{(k)}\!+\!\bm{\mu}_{d,b}^{(k)}\odot\frac{\partial{\tilde{R}_{b}^{(k)}}}{\partial{\bm{W}}_{d,b}}\right\}$;
@@
LAPGA implements an iterative optimizer characterized by a small number of iterations and low-complexity, approximated operations. By leveraging data-driven insights, it mitigates potential performance loss while significantly reducing computational overhead. Specifically, the average complexity per iteration is:
@@
which follows since the complexity in (21) is obtained by equal contributions of the digital and analog settings [22]. The reduction in (23) is shown in Section IV-D to have a significant impact on latency with minimal impact on performance.
IV-D Numerical Evaluation
In this section, we evaluate the performance of LAPGA through two numerical simulations. Our primary objective is to demonstrate that the proposed framework achieves performance comparable to non-approximated solvers while significantly reducing the online computational burden. To this end, we compare LAPGA against the following benchmarks, selected for their direct relationship to the PGA framework and their established efficacy in hybrid beamforming [3]:
@@ IV-D Numerical Evaluation
    Unfolded Full PGA: The learned architecture from [22] using exact gradients and trainable scalar step-sizes.
@@
    Classic PGA: The traditional iterative solver using manually optimized, fixed step-sizes.
We first consider a small-scale MIMO configuration with $B=8$ frequency bins, $N=6$ users, $L=10$ RF chains, and $M=12$ antennas. The channel realizations are generated using the QuadRiGa model [43], utilizing $1,000$ instances for training and $100$ for testing.
We implement LAPGA by employing the approximations in I1I3. Specifically, we set $\mathcal{K}_{a}^{\rm approx}=\{1,\dots,K\}$ and define $\mathcal{K}_{d,b}^{\rm approx}$ as the set of even iteration indices. This setup is further utilized to assess the individual impact of the proposed approximations. To this end, we compare LAPGA against the benchmarks B1B2 as well as the following architectural variations:
@@
Fig. 2 reports the average per-user sum-rate versus signal-to-noise ratio (SNR) (defined as $1/N\sigma^{2}$) for $K=5$ iterations, except for B2, which uses $K=50$. We observe in Fig. 2 that deep unfolding leads to significant performance gains. Specifically, B1 achieves better results than B2 while requiring $90\%$ fewer iterations, while LAPGA surpasses the rate obtained by B1 despite requiring $70\%$ fewer calculations. Overall, our proposed algorithm outperforms B2 while reducing the computational burden by $89\%$$97.3\%$, depending on the SNR.
To further quantify these gains, Table I summarizes the sum-rate (mean and standard deviation) at 0 dB and the total number of real-valued products required for the precoding task. The table highlights the dramatic complexity reduction of LAPGA. It also demonstrates that the approximations I1I2 allow for various balances between performance and complexity, and that the expanded parameterization in I3 is crucial for learning to mitigate the errors induced by these approximations.
Next, we evaluate the proposed framework in a larger MIMO setting, with $B=64$, $N=12$, $L=12$, and $M=32$. Building on the ablation study conducted in the previous section, we now consider LAPGA employing the full suite of approximations in I1I3. We maintain $K=5$ iterations for the unfolded optimizers, whereas B2 is configured with $K=100$ iterations to ensure convergence.
@@
The average per-user sum-rate versus SNR is illustrated in Fig. 3, while Table II provides a detailed complexity analysis at 0 dB. The results indicate that LAPGA achieves superior performance in both the low and high SNR regimes compared to the benchmarks. In the mid-range SNR, while B1 exhibits slightly higher rates, it entails a $70\%$ increase in computational overhead. Furthermore, B2 is shown to be three orders of magnitude more complex than our proposed approach.
@@ V Case Study 2: Robust PCA
Through the hybrid beamforming study, we highlighted how unfolding combined with approximated computations can enable efficient inference with drastically reduced latency. In this section, we explore the application of RPCA, which is a widely-used task in various fields ranging from background subtraction and anomaly detection in image and video processing [44]. RPCA serves as a representative setup for optimization problems that are $(i)$ computationally demanding and have motivated several unfolded, yet non-approximated, optimization strategies in prior work [27, 28]; $(ii)$ formulated using surrogate mathematical objectives rather than the true task loss. In particular, conventional RPCA formulations rely on convex relaxations of the underlying objective [30], which may be misaligned with the ultimate performance metric of interest. This makes RPCA an especially suitable domain for examining the added value of our approach in scenarios where the learning objective deviates from the optimization one.
To describe this case study, we first formulate RPCA and the considered first-order optimizer in Subsections V-A and V-B, respectively. Then we detail our unfolded methodology and evaluate it in Subsections V-C and V-D, respectively.